OpenStudy (anonymous):

Before a nuclear reaction, there is 0.0100 kg of uranium. After the nuclear reaction, there is 0.0090 kg of uranium. How much energy was released? (Points : 5) 5.0 × 10^-4 J 1.5 × 10^5 J 4.5 × 10^13 J 9.0 × 10^13 J

OpenStudy (michele_laino):

here we have to apply the Einstein's relationship: \[\Large E = \Delta m \times {c^2}\]

OpenStudy (michele_laino):

where \Delta m = 0.01-0.009=... and c=3*10^8 m/sec

OpenStudy (anonymous):

Ok thanks